A spring gun is set up to fire vertically. The spring is compressed a distance of 13 cm and has a spring constant of 410 N/m and the ball it fires has a mass of 180 grams. Determine the maximum height of the ball.
Answer #1

Solution) X = 13 cm = 0.13 m

K = 410 N/m

m = 180 g = 0.180 kg

H = ?

Equating potential energy = Spring potential energy

PE = U

mgH = (1/2)(K)(x^2)

0.180×9.8×H = (1/2)(410)(0.13^2)

0.180×9.8×H = 3.46

H = (3.46)/(0.180×9.8)

H = 1.96 m